OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. 1) y - 1 = 4(x - 3) 2) y + 1 = 4(x + 3) 3) y - 3 = 4(x + 1) 4) y + 3 = 4(x - 1)

OpenStudy (ankit042):

Y = mx+c, where m is the slope of the line.. use this to solve

OpenStudy (anonymous):

use point slope form..(y-y0)=m(x-x0)

OpenStudy (anonymous):

(y--1)=4(x--3) --->(y+1)=4(x+3) option 2